# Spectral results for free random variables

### v1

@@ Abstract.
$\varepsilon\mapsto\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)=\mathrm{tr}[\log((a-\lambda)^{\ast}(a-\lambda)+\varepsilon)^{-1}]$
admits a real analytic extension to a neighborhood of 0 in $\mathbb{R}.$ Then we will show that $\lambda$ is outside the spectrum of $a.$
@@ 1. Introduction
A paper of Gross and Malliavin [12] then gave a stochastic construction of the Segal–Bargmann over $K,$ using the Brownian motions in $K$ and $K_{\mathbb{C}}$ and methods from [11]. Finally, Biane [4] essentially took the construction of Gross and Malliavin for the case $K=U(N)$ and $K_{\mathbb{C}}=GL(N;\mathbb{C})$ and took the limit as $N\rightarrow\infty.$ Biane’s work indicated a close relationship between the free unitary Brownian motion (large-$N$ limit of Brownian motion in $U(N)$) and the free multiplicative Brownian motion (large-$N$ limit of Brownian motion in $GL(N;\mathbb{C})$).
Biane’s work then provided the motivation and technical tools for work of Hall–Kemp [21] computing the support of the Brown measure of the free multiplicative Brownian motion and then work of Driver–Hall–Kemp [8] computing the Brown measure itself. (Here “Brown measure” [5] is a construction in free probability that mimics the notion of eigenvalue distribution in random matrix theory.) The paper [8] introduced a new PDE method for computing Brown measures, which has then been used in subsequent works of Ho–Zhong [23], Hall–Ho [17, 18], Demni–Hamdi [7], and Eaknipitsari–Hall [9]. We also mention the work of Zhong [31], which does not use the PDE method but obtains similar formulas using free probability methods. The present paper also uses the PDE method in the applications of our general result.
@@
We now briefly summarize the applications we will make of Theorem 2.6. We consider a element of the form $x+c$, where $c$ is circular and $x$ is freely independent of $x,$ or more generally $x+g,$ where $g$ is elliptic and freely independent of $x.$ Assume that the spectrum of $x$ coincides with the support of its Brown measure, which will happen, for example, if $x$ is normal. Then the spectrum of $x+c$ coincides with the support of its Brown measure (Corollary 4.8), and the same result holds more generally for $x+g$ (Theorem 4.15). Meanwhile, consider the general free multiplicative Brownian motion $b_{s,\tau}$ introduced in [18, Section 2.1] and let $u$ be a unitary element that is freely independent of $b_{s,\tau}.$ Then the spectrum of $ub_{s,\tau}$ coincides with the support of its Brown measure (Theorem 5.8). In particular, we determine the spectrum of the free multiplicative Brownian motion $b_{t},$ in the original form introduced by Biane [4, Section 4.2.1]. The spectrum of $b_{t}$ is equal to the support of its Brown measure, which is the closure of the domain $\Sigma_{t}$ in [8]. See Figure 1.
@@ 2.1. Brown measure
for each Borel set $E,$ where $\nu_{a}$ is the projection-valued measure associated to $a$ by the spectral theorem (e.g., [15, Theorem 7.12]). The measure $\mu_{a}$ is the unique compactly supported probability measure on $\mathbb{C}$ satisfying
@@ 2.3. The regularized log potential and its derivative
(See Section 11.5 of the monograph [26] of Mingo and Speicher.) Then $S$ is a $C^{\infty}$ function of both $\lambda$ and $\varepsilon$ and is subharmonic [26, Equation (11.8)] as a function of $\lambda$ for fixed $\varepsilon.$ For $\lambda\in\mathbb{C}$ fixed, $S(\lambda,\varepsilon)$ decreases as $\varepsilon$ decreases, so that the limit as $\varepsilon\rightarrow 0^{+}$ exists, possibly equal to $\infty.$ After separating the log function into its positive and negative parts and applying monotone convergence, we find that
@@
whenever $a(\cdot)$ is a differentiable function with values in the space of positive elements of $\mathcal{A}.$ (See [5, Lemma 1.1] or [16, Equation (25)].) Using this result, we compute that
@@
See [31, Theorem 4.6]. This result is a strengthening of a result of Hall–Kemp [21, Theorem 1.2], which requires finiteness (and local boundedness) of the quantity $\mathrm{tr}[\left|(a-\lambda)^{2}\right|^{-2}].$
@@ 2.4. The main result
Meanwhile, it is known [24, 6.1.3 Proposition] that $u$ and $p$ must also belong to $\mathcal{A}.$ Then by the cyclic property of the trace,
@@
which is a holomorphic function of $\lambda\in\mathbb{C}\setminus[0,\infty).$ Note from (2.7) that
@@
$(\left|a-\lambda\right|^{2}+\varepsilon)^{-1}$ $=\left|a-\lambda\right|^{2}(1+\varepsilon\left|a-\lambda\right|^{-2})$
$=\left|a-\lambda\right|^{2}\sum_{k=0}^{\infty}(-1)^{k}\varepsilon^{k}\left|a-\lambda\right|^{-2k}.$
@@ 3.1. Additive case
and any pair $t>0$ and $\gamma\in\mathbb{C}$ satisfying (3.5) arises for some choice of $a,$ $b,$ and $\theta.$ (See [18, Section 2.1] or [31, Section 2.4].) We use the notation $g_{t,\gamma}$ to denote such an element.
@@
Ho and Zhong [23, Section 3] computed the Brown measure of an element of the form $x+c_{t},$ where $x$ is self-adjoint and freely independent of $c_{t}.$ See Figure 2. Zhong then computed the Brown measure of $x+c_{t},$ where $x$ is freely independent of $c_{t}$ but otherwise arbitrary. See Figure 3.
@@
In the circular case, $\mathrm{supp}(\mathrm{Br}_{x+c_{t}})$ was computed—by Ho–Zhong [23, Section 3] when $x$ is self-adjoint and by Zhong [31] when $x$ is arbitrary—as the closure $\overline{\Sigma}_{t}$ of a certain domain $\Sigma_{t}.$ Then by results of [31], $\mathrm{supp}(\mathrm{Br}_{x+g_{t,\gamma}})$ is the image of $\mathrm{supp}(\mathrm{Br}_{x+c_{t}})$ under a certain explicit map $\Phi_{t,\gamma}.$ See Section 4.3 for more information.
@@ 3.2. Multiplicative case
Hall and Kemp showed that the support of Brown measure of $b_{t}$ is contained in the closure of a certain set $\Sigma_{t}$, which was introduced by Biane in [4, Section 4.2.6]. Driver, Hall, and Kemp [8] then computed the Brown measure of $b_{t}$ and showed that its support is exactly $\overline{\Sigma}_{t}.$ Ho and Zhong [23, Section 4] extended the results of [8] to compute the Brown measure of $ub_{t},$ where $u$ is a unitary element that is freely independent of $b_{t}$. Finally, Hall and Ho computed the Brown measure of $ub_{t,\gamma}$ for general $t$ and $\gamma.$ See Figures 5 and 6.
@@
For more detailed statements, see Theorem 5.4 in the case of $ub_{t}$ and Theorem 5.8 in the case of $ub_{t,\gamma}.$ As in the additive case, the proof of the theorem relies on the prior computation of the Brown support, in [8] for the case of $b_{t}$ itself, in [23, Section 4] for $ub_{t},$ and in [18] for $ub_{t,\gamma}$.
@@
For a measure $\mu$ on $\mathbb{R},$ we let $\tilde{\mu}$ be the symmetrization of $\mu,$ that is, the average of $\mu$ and its push-forward under the map $\xi\mapsto-\xi.$ We first note that the symmetrization of $\mu_{|c_{t}|}$ is the semicircular law $\mathrm{sc}_{t}$ of variance $t$ [26, p. 174]. By Proposition 3.5 of [13], $x+c_{t}$ is $R$-diagonal and the law of $|x+c_{t}|$, determined by its symmetrization, is the free convolution
@@ 4.1. The self-adjoint plus circular case
Ho and Zhong [23, Section 3] compute the Brown measure of an element of the form $x+c_{t},$ where $x$ is self-adjoint, $c_{t}$ is circular of variance $t,$ and $x$ and $c_{t}$ are freely independent. We now introduce some of the PDE techniques from [23] that we will use to prove a result about the spectrum of $x+c_{t}.$ (See also [16] for a gentle introduction to the PDE method.) We consider the regularized log potential of $x+c_{t}$, as in (2.3), which we write as
@@
According to [23, Proposition 3.2], $S$ satisfies the PDE
@@
where $\mathrm{tr}[\log(\left|x-\lambda\right|^{2})]$ is well defined and harmonic for $\lambda$ outside $\bar{\Sigma}_{t},$ by Lemma 4.2. Thus, the Brown measure is zero outside $\overline{\Sigma}_{t}.$ There is a different analysis in [23, Section 3.2.2] to actually compute the Brown measure, inside $\Sigma_{t}$, but this does not concern us here—except for the result [23, Theorem 3.13] that the support of $\mathrm{Br}_{x+c_{t}}$ is equal to (not just contained in) $\overline{\Sigma}_{t}.$
@@
Let $c_{t}$ be circular of variance $t$, let $x$ be self-adjoint and freely independent of $c_{t},$ and consider the function $S$ in (4.1). Then for each $\lambda$ outside of $\overline{\Sigma}_{t},$ the function $\frac{\partial S}{\partial\varepsilon}(t,\lambda,\varepsilon)$ is analytic at $\varepsilon=0.$ Thus, by Theorem 2.6, the spectrum of $x+c_{t}$ is contained in $\overline{\Sigma}_{t}.$ Since [23, Theorem 3.13] tells us that the support of $\mathrm{Br}_{x+c_{t}}$ is exactly $\overline{\Sigma}_{t}$ (and since the Brown measure of any element is always supported on its spectrum), we conclude that the spectrum of $x+c_{t}$ coincides with its Brown support.
@@
The function $t_{\ast}(\lambda,\varepsilon_{0})$ in (4.12) is continuous in $\lambda$ for $\varepsilon_{0}>0$. As $\varepsilon_{0}$ decreases to 0, $t_{\ast}(\lambda,\varepsilon_{0})$ decreases to $T(\lambda).$ It then follows from an elementary result (e.g., [30, Theorem 15.84]) that $T$ is upper semicontinuous and therefore that $\Sigma_{t}$ is an open set. ∎
@@
Now, by the standard formula for the derivative of the inverse (e.g., [16, Equation (24)]), we have
@@ 4.2. Arbitrary plus circular
Since [31, Theorem B] tells us that the support of $\mathrm{Br}_{x+c_{t}}$ is exactly $\overline{\Sigma}_{t}$ (and since the Brown measure of any element is always supported on its spectrum), we conclude that
@@ 4.3. Arbitrary plus elliptic
We first let $\varepsilon\rightarrow 0$ in the second part of Eq. (3.33) in [31], which tells us that $p_{\lambda}^{(0)}(w(\varepsilon))=p_{\lambda}^{c,(t)}(\varepsilon),$ where these quantities are defined in Notation 3.10 of [31]. Now, for $\lambda$ outside $\overline{\Sigma}_{t},$ the quantity $w(\varepsilon)$ will tend to zero as $\varepsilon\rightarrow 0$ by [31, Lemma 3.5]. But by our assumptions, $\lambda\in(\overline{\Sigma}_{t})^{c}$ is also outside the spectrum of $x.$ Thus, the quantity $p_{\lambda}^{(0)}(w(\varepsilon))$ in Notation 3.10 in [31] will converge to
@@
This result is the first relation in [31, Corollary 3.11], where the relation between $z$ and $\lambda$ there is given by Equation (3.27) and the last displayed equation in the proof of Proposition 5.2. One may also obtain (4.28) from the Hamilton–Jacobi analysis of the PDE in (4.18) or (4.19). From that perspective, (4.28) amounts to second Hamilton–Jacobi formula (4.8), along with the fact that—since $\varepsilon$ does not appear explicitly on the right-hand side of (4.18) or (4.19)— $p_{\varepsilon}$ is a constant of motion.
@@
Now, the support of the Brown measure of $x+c_{t}$ is contained in $\overline{\Sigma}_{t}$ [23, Theorem 3.8], from which we can see that the map $\Phi_{t,\gamma}$ in (4.20) is holomorphic on $(\overline{\Sigma}_{t})^{c}.$ Thus, $(\Phi_{t,\gamma})_{\ast}$ is just the holomorphic derivative (a complex number), interpreted as a $2\times 2$ matrix. But by Lemma 4.14, $\Phi_{t,\gamma}$ is injective on $(\overline{\Sigma}_{t})^{c},$ which means that the holomorphic derivative can never vanish.
@@
The continuity of $\Phi_{t,\gamma}$ follows from Lemma 5.11 in [31]. We will then follow one of the standard proofs of the fundamental theorem of algebra, using the concept of the fundamental group (e.g., [27, Theorem 56.1]). The Cauchy transform $G_{x+c_{t}}(\lambda)$ in the definition (4.20) of $\Phi_{t,\gamma}(\lambda)$ behaves like $1/\lambda$ near infinity. Thus,

### v2

@@ Abstract.
$\varepsilon\mapsto\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)=\mathrm{tr}[((a-\lambda)^{\ast}(a-\lambda)+\varepsilon)^{-1}]$
admits a real-analytic extension to a neighborhood of 0 in $\mathbb{R}.$ Then we will show that $\lambda$ is outside the spectrum of $a.$
@@ 1. Introduction
A paper of Gross and Malliavin [14] then gave a stochastic construction of the Segal–Bargmann transform over $K,$ using the Brownian motions in $K$ and $K_{\mathbb{C}}$ and methods from [13]. Finally, Biane [4] essentially took the construction of Gross and Malliavin for the case $K=U(N)$ and $K_{\mathbb{C}}=GL(N;\mathbb{C})$ and took the limit as $N\rightarrow\infty.$ Biane’s work indicated a close relationship between the free unitary Brownian motion (large-$N$ limit of Brownian motion in $U(N)$) and the free multiplicative Brownian motion (large-$N$ limit of Brownian motion in $GL(N;\mathbb{C})$).
Biane’s work then provided the motivation and technical tools for work of Hall–Kemp [23] computing the support of the Brown measure of the free multiplicative Brownian motion and then work of Driver–Hall–Kemp [9] computing the Brown measure itself. (Here “Brown measure” [6] is a von Neumann algebra construction that mimics the notion of eigenvalue distribution in random matrix theory.) The paper [9] introduced a new PDE method for computing Brown measures, which has then been used in subsequent works of Ho–Zhong [25], Hall–Ho [19, 20], Demni–Hamdi [8], and Eaknipitsari–Hall [10]. We also mention the work of Zhong [34], which does not use the PDE method but obtains similar formulas using free probability methods. The present paper also uses the PDE method in the applications of our general result.
@@
We now briefly summarize the applications we will make of Theorem 2.6. We consider a element of the form $x+c$, where $c$ is circular and $x$ is freely independent of $x,$ or more generally $x+g,$ where $g$ is elliptic and freely independent of $x.$ Assume that the spectrum of $x$ coincides with the support of its Brown measure, which will happen, for example, if $x$ is normal. Then the spectrum of $x+c$ coincides with the support of its Brown measure (Corollary 4.8), and the same result holds more generally for $x+g$ (Theorem 4.15). Meanwhile, consider the general free multiplicative Brownian motion $b_{s,\tau}$ introduced in [20, Section 2.1] and let $u$ be a unitary element that is freely independent of $b_{s,\tau}.$ Then the spectrum of $ub_{s,\tau}$ coincides with the support of its Brown measure (Theorem 5.8). In particular, we determine the spectrum of the free multiplicative Brownian motion $b_{t},$ in the original form introduced by Biane [4, Section 4.2.1]. The spectrum of $b_{t}$ is equal to the support of its Brown measure, which is the closure of the domain $\Sigma_{t}$ in [9]. See Figure 1.
@@ 2.1. Brown measure
for each Borel set $E,$ where $\nu_{a}$ is the projection-valued measure associated to $a$ by the spectral theorem (e.g., [17, Theorem 7.12]). The measure $\mu_{a}$ is the unique compactly supported probability measure on $\mathbb{C}$ satisfying
@@ 2.3. The regularized log potential and its derivative
(See Section 11.5 of the monograph [28] of Mingo and Speicher.) Then $S$ is a $C^{\infty}$ function of both $\lambda$ and $\varepsilon$ and is subharmonic [28, Equation (11.8)] as a function of $\lambda$ for fixed $\varepsilon.$ For $\lambda\in\mathbb{C}$ fixed, $S(\lambda,\varepsilon)$ decreases as $\varepsilon$ decreases, so that the limit as $\varepsilon\rightarrow 0^{+}$ exists, possibly equal to $\infty.$ After separating the log function into its positive and negative parts and applying monotone convergence, we find that
@@
whenever $a(\cdot)$ is a differentiable function with values in the space of positive elements of $\mathcal{A}.$ (See [6, Lemma 1.1] or [18, Equation (25)].) Using this result, we compute that
@@
See [34, Theorem 4.6]. This result is a strengthening of a result of Hall–Kemp [23, Theorem 1.2], which requires finiteness (and local boundedness) of the quantity $\mathrm{tr}[\left|(a-\lambda)^{2}\right|^{-2}].$
@@ 2.4. The main result
Meanwhile, it is known [26, 6.1.3 Proposition] that $u$ and $p$ must also belong to $\mathcal{A}.$ Then by the cyclic property of the trace,
@@
which is a holomorphic function of $z\in\mathbb{C}\setminus[0,\infty).$ Note from (2.7) that
@@
$(\left|a-\lambda\right|^{2}+\varepsilon)^{-1}$ $=\left|a-\lambda\right|^{-2}(1+\varepsilon\left|a-\lambda\right|^{-2})$
$=\left|a-\lambda\right|^{-2}\sum_{k=0}^{\infty}(-1)^{k}\varepsilon^{k}\left|a-\lambda\right|^{-2k}.$
@@ 3.1. Additive case
and any pair $t>0$ and $\gamma\in\mathbb{C}$ satisfying (3.5) arises for some choice of $a,$ $b,$ and $\theta.$ (See [20, Section 2.1] or [34, Section 2.4].) We use the notation $g_{t,\gamma}$ to denote such an element.
@@
Ho and Zhong [25, Section 3] computed the Brown measure of an element of the form $x+c_{t},$ where $x$ is self-adjoint and freely independent of $c_{t},$ building on earlier results of Biane and Lehner [5, Section 3]. See Figure 2. Zhong then computed the Brown measure of $x+c_{t},$ where $x$ is freely independent of $c_{t}$ but otherwise arbitrary. See Figure 3.
@@
In the circular case, $\mathrm{supp}(\mathrm{Br}_{x+c_{t}})$ was computed—by Ho–Zhong [25, Section 3] when $x$ is self-adjoint and by Zhong [34] when $x$ is arbitrary—as the closure $\overline{\Sigma}_{t}$ of a certain domain $\Sigma_{t}.$ Then by results of [34], $\mathrm{supp}(\mathrm{Br}_{x+g_{t,\gamma}})$ is the image of $\mathrm{supp}(\mathrm{Br}_{x+c_{t}})$ under a certain explicit map $\Phi_{t,\gamma}.$ See Section 4.3 for more information.
Further information about the domain $\Sigma_{t}$ in the preceding remark was obtained by Erdős and Ji in [11].
@@ 3.2. Multiplicative case
Hall and Kemp showed that the support of Brown measure of $b_{t}$ is contained in the closure of a certain set $\Sigma_{t}$, which was introduced by Biane in [4, Section 4.2.6]. Driver, Hall, and Kemp [9] then computed the Brown measure of $b_{t}$ and showed that its support is exactly $\overline{\Sigma}_{t}.$ Ho and Zhong [25, Section 4] extended the results of [9] to compute the Brown measure of $ub_{t},$ where $u$ is a unitary element that is freely independent of $b_{t}$. Finally, Hall and Ho computed the Brown measure of $ub_{t,\gamma}$ for general $t$ and $\gamma.$ See Figures 5 and 6.
@@
For more detailed statements, see Theorem 5.4 in the case of $ub_{t}$ and Theorem 5.8 in the case of $ub_{t,\gamma}.$ As in the additive case, the proof of the theorem relies on the prior computation of the Brown support, in [9] for the case of $b_{t}$ itself, in [25, Section 4] for $ub_{t},$ and in [20] for $ub_{t,\gamma}$.
@@
For a measure $\mu$ on $\mathbb{R},$ we let $\tilde{\mu}$ be the symmetrization of $\mu,$ that is, the average of $\mu$ and its push-forward under the map $\xi\mapsto-\xi.$ We first note that the symmetrization of $\mu_{|c_{t}|}$ is the semicircular law $\mathrm{sc}_{t}$ of variance $t$ [28, p. 174]. By Proposition 3.5 of [15], $x+c_{t}$ is $R$-diagonal and the law of $|x+c_{t}|$, determined by its symmetrization, is the free convolution
@@ 4.1. The self-adjoint plus circular case
Ho and Zhong [25, Section 3] compute the Brown measure of an element of the form $x+c_{t},$ where $x$ is self-adjoint, $c_{t}$ is circular of variance $t,$ and $x$ and $c_{t}$ are freely independent. We now introduce some of the PDE techniques from [25] that we will use to prove a result about the spectrum of $x+c_{t}.$ (See also [18] for a gentle introduction to the PDE method.) We consider the regularized log potential of $x+c_{t}$, as in (2.3), which we write as
@@
According to [25, Proposition 3.2], $S$ satisfies the PDE
@@
where $\mathrm{tr}[\log(\left|x-\lambda\right|^{2})]$ is well defined and harmonic for $\lambda$ outside $\bar{\Sigma}_{t},$ by Lemma 4.2. Thus, the Brown measure is zero outside $\overline{\Sigma}_{t}.$ There is a different analysis in [25, Section 3.2.2] to actually compute the Brown measure, inside $\Sigma_{t}$, but this does not concern us here—except for the result [25, Theorem 3.13] that the support of $\mathrm{Br}_{x+c_{t}}$ is equal to (not just contained in) $\overline{\Sigma}_{t}.$
@@
Let $c_{t}$ be circular of variance $t$, let $x$ be self-adjoint and freely independent of $c_{t},$ and consider the function $S$ in (4.1). Then for each $\lambda$ outside of $\overline{\Sigma}_{t},$ the function $\frac{\partial S}{\partial\varepsilon}(t,\lambda,\varepsilon)$ is analytic at $\varepsilon=0.$ Thus, by Theorem 2.6, the spectrum of $x+c_{t}$ is contained in $\overline{\Sigma}_{t}.$ Since [25, Theorem 3.13] tells us that the support of $\mathrm{Br}_{x+c_{t}}$ is exactly $\overline{\Sigma}_{t}$ (and since the Brown measure of any element is always supported on its spectrum), we conclude that the spectrum of $x+c_{t}$ coincides with its Brown support.
@@
The function $t_{\ast}(\lambda,\varepsilon_{0})$ in (4.12) is continuous in $\lambda$ for $\varepsilon_{0}>0$. As $\varepsilon_{0}$ decreases to 0, $t_{\ast}(\lambda,\varepsilon_{0})$ decreases to $T(\lambda).$ It then follows from an elementary result (e.g., [32, Theorem 15.84]) that $T$ is upper semicontinuous and therefore that $\Sigma_{t}$ is an open set. ∎
@@
Now, by the standard formula for the derivative of the inverse (e.g., [18, Equation (24)]), we have
@@ 4.2. Arbitrary plus circular
Since [34, Theorem B] tells us that the support of $\mathrm{Br}_{x+c_{t}}$ is exactly $\overline{\Sigma}_{t}$ (and since the Brown measure of any element is always supported on its spectrum), we conclude that
@@ 4.3. Arbitrary plus elliptic
We first let $\varepsilon\rightarrow 0$ in the second part of Eq. (3.33) in [34], which tells us that $p_{\lambda}^{(0)}(w(\varepsilon))=p_{\lambda}^{c,(t)}(\varepsilon),$ where these quantities are defined in Notation 3.10 of [34]. Now, for $\lambda$ outside $\overline{\Sigma}_{t},$ the quantity $w(\varepsilon)$ will tend to zero as $\varepsilon\rightarrow 0$ by [34, Lemma 3.5]. But by our assumptions, $\lambda\in(\overline{\Sigma}_{t})^{c}$ is also outside the spectrum of $x.$ Thus, the quantity $p_{\lambda}^{(0)}(w(\varepsilon))$ in Notation 3.10 in [34] will converge to
@@
This result is the first relation in [34, Corollary 3.11], where the relation between $z$ and $\lambda$ there is given by Equation (3.27) and the last displayed equation in the proof of Proposition 5.2. One may also obtain (4.28) from the Hamilton–Jacobi analysis of the PDE in (4.18) or (4.19). From that perspective, (4.28) amounts to second Hamilton–Jacobi formula (4.8), along with the fact that—since $\varepsilon$ does not appear explicitly on the right-hand side of (4.18) or (4.19)— $p_{\varepsilon}$ is a constant of motion.
@@
Now, the support of the Brown measure of $x+c_{t}$ is contained in $\overline{\Sigma}_{t}$ [25, Theorem 3.8], from which we can see that the map $\Phi_{t,\gamma}$ in (4.20) is holomorphic on $(\overline{\Sigma}_{t})^{c}.$ Thus, $(\Phi_{t,\gamma})_{\ast}$ is just the holomorphic derivative (a complex number), interpreted as a $2\times 2$ matrix. But by Lemma 4.14, $\Phi_{t,\gamma}$ is injective on $(\overline{\Sigma}_{t})^{c},$ which means that the holomorphic derivative can never vanish.
@@
The continuity of $\Phi_{t,\gamma}$ follows from Lemma 5.11 in [34]. We will then follow one of the standard proofs of the fundamental theorem of algebra, using the concept of the fundamental group (e.g., [29, Theorem 56.1]). The Cauchy transform $G_{x+c_{t}}(\lambda)$ in the definition (4.20) of $\Phi_{t,\gamma}(\lambda)$ behaves like $1/\lambda$ near infinity. Thus,